NDOS CommissionMtng 20121029¶
Power Outage on previous Monday -
It took days to recover. Mounts took awhile to recover. db-01 had to be reset
data disk machine came up checking dicks. We should have it check its disks on maintainence days.
Leon asks about auto mount.
What do we do on maintenance days?
Another 15 or so to go on detector.
APDs for the test rig are out for coating. (Missing spacer frames)
APD Test: Asked Jao to configure. Peter will ping to check status.
(DSO we've seen jumps correlated with the wrong multiplexing cycle.
Junting wants a temperature scan to look for thermal noise.
Voltage scan ?
Jan - DCS table was successfully copied. FEB shut off is working again. 14204-206
-14217 low slice count
-2-2 33 silicon sit off in subrun 2 over weekend